the measure of one complementary angles is fifteen less than twice the measure of the other.Find the measure of each angle
Answer by cleomenius(959)

x + 15 = 2(90-x)
x + 15 = 180 - 2x
3x = 165
x = 55 is one angle. the complement is 35
35 * 2 = 70
70 - 15 = 55 degrees.
Cleomenius.